OpenStudy (anonymous):

I am trying to get the following function in vertex form I need steps to get me there, can you help? The function is y=x^2+14x+55

OpenStudy (anonymous):

can you go through the whole process of completing the square? I have other problems like this and it will be easier for me to get those so i can use this one as an example

OpenStudy (campbell_st):

ok.. so the vertex form of a parabola, with a vertex at ( h, k) is \[y = ( x -h)^2 + k\] so in your question you need to form a perfect square and have constant term so if you look at your equation \[y = x^2 + 14x + 55\] you will use \[x^2 + 14x \] to form the perfect square... does that make sense..?

OpenStudy (anonymous):

yeah thats good thank you

OpenStudy (campbell_st):

ok... so to complete the square we need to find half of the coefficient of the middle term in general you have \[y = ax^2 + bx + c\] so start by finding b/2 in your question b = 14 so find half of it..

OpenStudy (anonymous):

well ultimately i am trying to find the vertex of a parabola so i think the equation is (y=a(x-k)^2+c where a =positive or negative 1 right? thats the equation i need more help with